OpenStudy (yanasidlinskiy):

Suppose you just received a shipment of twelve televisions. Two of the televisions are defective. If two televisions are randomly​ selected, compute the probability that both televisions work. What is the probability at least one of the two televisions does not​ work?

OpenStudy (yanasidlinskiy):

@ganeshie8

OpenStudy (yanasidlinskiy):

I keep getting stupid answers and it's really irritating me.

ganeshie8 (ganeshie8):

We are given that 2 tvs are defect out of 12 tvs. This means, P(good) = 10/12 P(bad) = 2/12

OpenStudy (yanasidlinskiy):

So.... I'd divide 2/12 right?

ganeshie8 (ganeshie8):

Lets do it like this

ganeshie8 (ganeshie8):

You pick one tv first. Whats the probability for it to be good ?

ganeshie8 (ganeshie8):

Number of good tvs = 10 Total number of tvs = 12 P(good) = 10/12 right?

ganeshie8 (ganeshie8):

After picking one tv, there will be 11 more tvs. Out of them, 9 tvs are good and 2 tvs are bad. So for picking a good tv the second time also, the probability would be : P("good second time") = 9/11 yes ?

ganeshie8 (ganeshie8):

Therefore, P("good both times") = 10/12 * 9/11 = 15/22

ganeshie8 (ganeshie8):

see if that makes any sense...

ganeshie8 (ganeshie8):

see if that makes any sense...

OpenStudy (yanasidlinskiy):

Okay, yes that does make sense. I was doing 2/12 and other crazy fractions that I don't know... But yes, that definitely makes sense. Great explanation!! Thanks. Switch brains with me please lol

ganeshie8 (ganeshie8):

Haha ikr, probability is a bit confusing in the start... But it helps if we take time to understand the problem thoroughly before attempting the solution..

ganeshie8 (ganeshie8):

Look at second part of the question

ganeshie8 (ganeshie8):

`What is the probability at least one of the two televisions does not​ work?`

OpenStudy (yanasidlinskiy):

So wait, it would be 15/22?

ganeshie8 (ganeshie8):

`. If two televisions are randomly​ selected, compute the probability that both televisions work.` 15/22 is the answer to this part of the problem

ganeshie8 (ganeshie8):

we still need to work the second part of the problem

OpenStudy (yanasidlinskiy):

I think I know this one.

OpenStudy (yanasidlinskiy):

I'd have: 1-0.682 Right?

OpenStudy (yanasidlinskiy):

Nevermind I got it.